if a generator is stepping up a 47.20 v to 75,000 V for interstate and he has a generated current of 4.258. how much current will be added to the interstate transmission line ?
Answer #1

here,

the voltage of primary , Vp = 47.2 V

the voltage of seccondary , Vs = 75000 V

the currnt in primary , Ip = 4.258 A

let the current in the seccondary be Is

the power remains the same

Vp * Ip = Vs * Is

47.2 * 4.258 = 75000 * Is

solving for Is

Is = 2.68 * 10^-3 A = 2.68 mA